OpenStudy (anonymous):

Is this correct ? Express the radical in simplest form: sqrt 40 would be 2 sqrt 20 or 4 sqrt 10

OpenStudy (anonymous):

\[\sqrt{40} = 2\sqrt{20}\]

OpenStudy (lgbasallote):

try it like this.. sqrt 40 = sqrt (4 x 10) = sqrt (4) x sqrt (10) do you get the answer now? :D

OpenStudy (lgbasallote):

if you do..pls tell the answer so i may know if you get it ^_^

OpenStudy (anonymous):

\[4\sqrt{10}\]

OpenStudy (lgbasallote):

no...what is square root of 4?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

\[2\sqrt{10}\]

OpenStudy (lgbasallote):

YES!!! yayyy :DDD

OpenStudy (anonymous):

thank you (:

OpenStudy (lgbasallote):

<tips imaginary hat>